# Optimizing importance weighting in the presence of sub-population shifts

### v1

@@ B.4 Numerical comparison
We can observe again in this figure that as $n$ decreases or $d$ increases, the variance becomes a greater share of the expected loss. This variance can be reduced by selecting a $p$ closer to the original $p_{\mathrm{tr}}$. However, this this subsequently leads to an increase in the bias. The optimal $p$ provides a trade-off between these two objectives.
@@ E.1 Datasets
An overview of the datasets, their group definitions and sizes, as well as the train/validation split used in the experiments, can be found in Table 4.
@@
Waterbirds: this dataset from Sagawa et al. (2020a) is a combination of the Places dataset (Zhou et al., 2016) and the CUB dataset (Welinder et al., 2010). A ‘water background’ is set by selecting an image from the lake and ocean categories in the places dataset, and the ‘land background’ is set based on the broadleaf and bamboo forest categories. A waterbird/land is then pasted in front of the background. The goal is to predict whether or not a picture contains a waterbird or landbird.
@@ E.3 Implementation details of existing methods
For each run, the hyperparameters are selected based on the worst-group accuracy on the respective validation set. For the L1 penalty, we select it from the following values: $0.1,1.0,3.3,10.0,33.33,100,300,500$. In the case that an multiple L1 penalties lead to the same worst-group accuracy, we select the highest one.
@@ E.4 Implementation details of optimizing the weights
Following the notation in Algorithm 1. We use a learning rate $\eta$ of 0.1 for each dataset and method, except in two cases: optimizing the weights for DFR on Waterbirds and MultiNLI, where we use a learning rate of 0.01 and 0.2 respectively. This was done after observing that the optimization procedure did not converge to a lower validation loss for the original choice of the learning rate. In the case of Waterbirds, we run Algorithm 1 for $T=200$ steps, and for $T=100$ in the case of CelebA and MultiNLI.

### v2

@@ B.4 Numerical comparison
We can observe in this figure that as $n$ decreases or $d$ increases, the variance becomes a greater share of the expected loss. This variance can be reduced by selecting a $p$ closer to the original $p_{\mathrm{tr}}$. However, this subsequently leads to an increase in the bias. The optimal $p$ provides a trade-off between these two objectives.
@@ E.1 Datasets
An overview of the datasets, their group definitions and sizes, as well as the train/validation split used in the experiments, is provided in Table 4.
@@
Waterbirds: this dataset from Sagawa et al. (2020a) is a combination of the Places dataset (Zhou et al., 2016) and the CUB dataset (Welinder et al., 2010). A ‘water background’ is set by selecting an image from the lake and ocean categories in the places dataset, and the ‘land background’ is set based on the broadleaf and bamboo forest categories. A waterbird/landbird is then pasted in front of the background. The goal is to predict whether or not a picture contains a waterbird or landbird.
@@ E.3 Implementation details of existing methods
For each run, the hyperparameters are selected based on the worst-group accuracy on the respective validation set. For the L1 penalty, we select it from the following values: $0.1,1.0,3.3,10.0,33.33,100,300,500$. In the case that multiple L1 penalties lead to the same worst-group accuracy, we select the highest one.
@@ E.4 Implementation details of optimizing the weights
Following the notation in Algorithm 1, we use a learning rate $\eta$ of 0.1 for each dataset and method, except in two cases: optimizing the weights for DFR on Waterbirds and MultiNLI, where we use a learning rate of 0.01 and 0.2 respectively. This was done after observing that the optimization procedure did not converge to a lower validation loss for the original choice of the learning rate. In the case of Waterbirds, we run Algorithm 1 for $T=200$ steps, and for $T=100$ in the case of CelebA and MultiNLI.